OpenStudy (anonymous):

Two number cubes are rolled --one blue and one yellow. Find the probability that the yellow cube shows a multiple of 3, given that the product is 6.

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (fibonaccichick666):

tell me what you think

OpenStudy (anonymous):

well there is 1 and 3 only so 2/6

OpenStudy (fibonaccichick666):

wait,

OpenStudy (fibonaccichick666):

first, how many ways can you get a product of 6?

OpenStudy (fibonaccichick666):

name them

OpenStudy (anonymous):

1x6 2x3 3x2 6x1

OpenStudy (fibonaccichick666):

good now lwt's say the yelow cube is first, how many are divisible by 3?

OpenStudy (fibonaccichick666):

and which are they

OpenStudy (anonymous):

2 of them, 6 and 3

OpenStudy (fibonaccichick666):

good so now,what is the probability of the occurrence of a multiple of three?

OpenStudy (anonymous):

2/6

OpenStudy (fibonaccichick666):

now, for the product to be six, how many ways can the second die go?

OpenStudy (anonymous):

it can roll a 2 and a 1

OpenStudy (fibonaccichick666):

so, that probability is?

OpenStudy (anonymous):

2/6

OpenStudy (anonymous):

so 4/36?

OpenStudy (anonymous):

1/9?

OpenStudy (anonymous):

wait no thats wrong

OpenStudy (fibonaccichick666):

hm?

OpenStudy (anonymous):

P(B|A) = 1/6 P(B|A) = 1/8 P(B|A) = 1/4 P(B|A) = 1/2

OpenStudy (anonymous):

those are my options

OpenStudy (fibonaccichick666):

how many ways are there to get a product of 6?

OpenStudy (anonymous):

4

OpenStudy (fibonaccichick666):

we are given that the product must be 6

OpenStudy (fibonaccichick666):

so if there are 4 ways to get 6

OpenStudy (anonymous):

the possible multiples for cube A are 3 and 6 therefore the probability is 2/6/. The product must be 6 therefore the second die has to roll a 2 or a 1 only

OpenStudy (anonymous):

therefore2/6*2/6-4/36

OpenStudy (fibonaccichick666):

yea, I agree... hmmm

OpenStudy (anonymous):

if 4/36 is simplified then that mean 2/18 and then 1/9

OpenStudy (anonymous):

this homework is a monster!!!!!!!!!!!!!

OpenStudy (fibonaccichick666):

i think it should be something like 4/36*2/6*2/6 but I'm not sure

OpenStudy (fibonaccichick666):

no non no so 2/6*2/4

OpenStudy (anonymous):

1/6

OpenStudy (fibonaccichick666):

yea